OpenStudy (love_333):

Medal + fan: Concrete can be bought by the cubic yard. How much will it cost to pour a slab 18 feet by 18 feet by 6 in for a patio if the concrete costs $42.00 per cubic yard? A $252 B $756 C $1,134 D $6,804

OpenStudy (love_333):

@Erak @satellite73

OpenStudy (erak):

Oh god, imperial unit conversions. So. 12 in = 1 ft, and 36 in = 1 yard, turn them all into yard measurements

OpenStudy (love_333):

18ft = 6yd 18ft=6yd 6in=1/2ft=.166yd

OpenStudy (love_333):

Did I do that right? lol

OpenStudy (erak):

okay the last one you can leave it was like (1/6)yd. Now do V = 6x6x(1/6), then multiply by the price

OpenStudy (love_333):

I got 252! a. :)

OpenStudy (erak):

yeah

OpenStudy (mathmale):

Be sure to use $ as your unit of measurement!

OpenStudy (love_333):

Thx for the reminder! @mathmale